becky has a board that is 2(3/4)feet long. she needs to cut the board into 11 peices of equal length. how long should each piece be?
Answer by hovuquocan1997(83)

You just need to take the total length and divide it by 11 :)
2 3/4 = 11/4
11/4 : 11 = 1/4
So each piece should be 1/4 feet long
TA-DAH
:D